# Learning an Ensemble Token from Task-driven Priors in Facial Analysis

### v1

@@ Abstract
Facial analysis exhibits task-specific feature variations. While Convolutional Neural Networks (CNNs) have enabled the fine-grained representation of spatial information, Vision Transformers (ViTs) have facilitated the representation of semantic information at the patch level. Although the generalization of conventional methodologies has advanced visual interpretability, there remains paucity of research that preserves the unified feature representation on single task learning during the training process. In this work, we introduce ET-Fuser, a novel methodology for learning ensemble token by leveraging attention mechanisms based on task priors derived from pre-trained models for facial analysis. Specifically, we propose a robust prior unification learning method that generates a ensemble token within a self-attention mechanism, which shares the mutual information along the pre-trained encoders. This ensemble token approach offers high efficiency with negligible computational cost. Our results show improvements across a variety of facial analysis, with statistically significant enhancements observed in the feature representations.
@@ 1 Introduction
Facial analysis exhibits feature variations depending on the task. Conversely, a unified feature representation derived from diverse features demonstrates robust performance across facial analysis tasks. Through neural network on deep learning, spatial and semantic features are represented from a single facial image. The Convolutional Neural Networks (CNNs) enabled the representation of spatial information, and more recently, the Vision Transformers (ViTs) has facilitated the representation of semantic information at the patch level. Notably, the attention mechanism within the transformer demonstrates robust representation even for inputs scattered in disparate spaces or across multi-modality [15, 43].
@@
Figure 1: The key concept of our proposed approach. Our method leverages pre-trained priors from a frozen encoder and unifies them into a learnable token, called ensemble token.
@@
The emergence of general models such as DINOv2 [27] and SAM [16] capable of handling multiple tasks has significantly advanced interpretability within the computer vision community. These general models have demonstrated significant achievements in image classification, segmentation, and detection, and as well as handle multi-tasks within a single encoder. Furthermore, linear probing of pre-trained models with unified feature representation has shown promising results [12, 18, 29, 14]. In the areas of multi-modality, CLIP [31, 35] and attention mechanisms [41, 6] has demonstrated promising results, confirming that unified feature representation significantly enhances model’s interpretability.
Compared to CNNs, ViTs can be less effective when addressing tasks constrained by specific domains or limited data. In fields on facial analysis, where data availability is often a limiting factor and the representation of fine-grained spatial details is crucial, these limitations become particularly pronounced. To mitigate these troublesome, the hybrid architecture or an ensemble of CNNs and ViTs in embedding space have recently proved prominence in the recent research [33, 42, 45]. In facial analysis tasks that emphasize spatial representation such as facial landmark detection and age estimation, CNNs serve as the primary encoder for learning canonical representations in spatial information. To further capture semantic information, self-attention mechanisms are subsequently applied to the spatially-informed feature representations. To effectively integrate both spatial and semantic information, a dual-path architecture [7, 47] is employed by featuring a CNN-based encoder alongside a ViT-based encoder. This architecture enables the fusion of spatial and semantic information at the feature level, resulting in a unified visual token.
@@
As noted, the capacity of single and multi task learning in conventional methodologies has stepped up the visual interpretability. However, there remains paucity of research that preserves the unified feature representation or encoder’s priors on single task learning during the training process. Although task-oriented fine-tuning can be achieved through ensemble learning [10, 23, 24], there is still a challenge in that the computational cost escalates proportionally to both the number of tasks and the encoder model’s capacity. In this research, we introduce a methodology for learning ensemble tokens by leveraging the attention mechanisms based on the task priors of pre-trained models in facial analysis tasks. The principal contributions of this study are shown in Fig. 1 and as follows:
  • We proposes ET-Fuser, a ensemble token fuser for robust prior unification learning based on self-attention mechanism, which shares the mutual information along the prior tasks.
@@
  • The ensemble token within the ET-Fuser denotes a highly efficient method and requires negligible computational cost. Furthermore, the ET-Fuser’s flexible architecture allows for straightforward adaptation and integration across domains and tasks
@@
  • We present a efficient task-specific adaptation without training from scratch, thereby remaining unaffected by constraints related to data limitations availability and hardware cost.
@@ 2 Related works
Figure 2: Ensemble learning typically train (red) pre-trained (blue) priors within a fusion module, while ControlNet incorporates a auxiliary light-weight (dashed-line) encoder. In contrast, ET-Fuser implicitly (grey) learns encoder priors within its fusion module, thereby reducing computational cost dramatically.
@@ 2.1 Leveraging pre-trained priors
Ensemble learning. The capacity to train specialized models for individual tasks has enabled ensemble methods to achieve robust performance through the exploitation of task-specific priors. These methods encompass a wide spectrum of strategies, ranging from classical techniques like bagging, boosting, and stacking, to more sophisticated fusion approaches such as majority voting and deep learning-based methodologies [30, 3, 19, 22]. Nevertheless, a fundamental limitation of ensemble methodologies resides in their computational cost, which is directly proportional to both model capacity and ensemble cardinality. This computational cost can preclude their applicability in real-time facial analysis scenarios as shown in Fig. 2 .
Knowledge distillation (KD). Leveraging the priors of pre-trained and progressive teacher model, KD constitutes an indispensable methodology in facial analysis, particularly in scenarios necessitating smaller model capacity. Yu et al. [48] introduced ICD-Face, which employs feature banks to provide sufficient high-quality positive pairs and incorporates a similarity distribution consistency loss between the probability distributions of teacher and student models. AdaDistill [1] further integrates KD within the context of face recognition, training the student model via a margin penalty softmax function. While KD facilitates performance enhancements not only through knowledge transfer from larger capacity models but also via self-distillation [52] within models of comparable capacity, training challenges may manifest when addressing tasks characterized by divergent embedding spaces or substantial disparities in teacher model’s capacity.
Low rank adaptation (LoRA). The introduction of LoRA [13] was driven by the need for parameter-efficient fine-tuning (PEFT), primarily to mitigate the computational cost and substantial hardware demands inherent in training large language models. Within the computer vision community, LoRA has also proven valuable as a PEFT strategy, particularly for diffusion models [11, 9]. Subsequent to the LoRA, ControlNet-based [49] adaptation of foundation models has demonstrated significant efficacy in domain and task adaptation even on multi-modality in face analysis [17, 39]. While ControlNet has established its capacity to preserve pre-trained priors within task-specific models, the potential for ensembling multiple encoders to achieve controllable training warrants further investigation.
@@ 2.2 Feature tokenization on facial analysis
The attention mechanism [6] and CLIP [28] have become a cornerstones within the computer vision community, effectively capturing both the semantic richness of individual tokens and the spatial hierarchies within single and multi-modality. In facial analysis, the HOTformer [37] introduced a novel approach employing lightweight CNNs guided by holistic and atomic tokens. This hybrid token of HOTformer effectively integrates facial semantics and outperforms the accuracy on face recognition. The Zhang et al. [50] proposed TokenFace, a transformer-based 3D face reconstruction model with a learnable facial component tokens encompassing shape, expressions, jaw pose, camera pose, texture, and lighting. This multi-modality effectively integrates the interdependencies between facial characteristics and image token embeddings derived from ViT within the single task. The POSTER++ [21] advanced the state-of-the-art in facial emotion recognition by replacing traditional pyramid architectures with a more efficient design, achieving performance with minimal computational cost. This method leverages cross-attention to fuse landmark priors and image features. The FaceXFormer [26] fuses the hierarchical representation extracted from CNN blocks. Its task-specific tokenization scheme facilitates multi-task learning within a single encoder architecture, effectively utilizing both self-attention and cross-attention mechanisms.
While attention mechanisms have demonstrated efficacy in capturing mutual information at both coarse and fine-grained levels, and while facial analysis benefits from the flexibility of determining between CNN- and ViT-based architecture contingent on task-specific requirements, current methodologies employ a single learnable encoder for mutual information learning. This architectural characteristic can induce performance bottlenecks, particularly when constrained by encoder’s capacity. So as in leveraging the pre-trained priors, ensemble learning often achieve state-of-the-art results, but their computational expense can preclude their practical application on facial analysis. Although LoRA-based architectures transform embedding representations via auxiliary layers, the use of a single, lightweight encoder prove insufficient to fully capture the overfull information encoded within each pre-trained priors. Our ET-Fuser addresses both the computational cost and the issue of prior’s redundancies through the ensemble token.
@@
Figure 3: Training phrase of proposed method. Within the designated blue block, all encoders $E$ are maintained as frozen layers, indicating that they are not subject to parameter updates during training. Both the ET-Fuser and the canonical branch decoder $d_{c}$ represent the trainable layer within the architecture. In the green block, decoders belonging to the set $D\setminus\{d_{c}\}$, which correspond to other tasks distinct from the canonical task, are excluded from the forward propagation graph. Furthermore, The existence of the ground truth set $Y\setminus\{y_{c}\}$, denoted as dashed line, is contingent upon the dataset, exhibiting potential variability in its presence; exceptionally, the $y_{c}$ associated with the canonical branch is an indispensable element.
@@ 3 Methodology
Inspired by the object token in DETR [2] and mask token in MaskFormer [4, 5], we propose learnable ensemble tokens corresponding to learnable query features in preliminaries. DETR introduced an anchor-free semantic representation leveraging object tokens. Subsequently, Mask2Former utilized masked attention to incorporate both semantic and spatial information into learnable query features. Motivated by the concept of learnable query embeddings, we introduce a ensemble token designed to capture diverse embeddings derived from the priors of a pre-trained model. While conceptually similar to a module that fuses information acquired from an ensemble, this ensemble token is specifically tailored to the canonical task, selectively attending to and learning pertinent semantic information.
3.1 Acquiring prior tokens from encoders
@@
In the preliminary stage, a set of N pre-trained encoders $E=\{\varepsilon_{1},\varepsilon_{2},...,\varepsilon_{n-1},\varepsilon_{c}\}$ corresponding to the expertise of each task is prepared. These encoders could be employed either CNN or ViT architectures, or even a hybrid thereof dictated by task-specific priors. Subsequent to the preparation of $E$, a canonical task branch is selected for task-specific training. As illustrated in Fig. 3, face landmark detection serves as the canonical task in the presented example. During training stage, supervision learning is exclusively derived from the ground truth $Y=\{y_{1},y_{2},...,y_{n-1},y_{c}\}$ corresponding to the designated canonical task; ground truth of the N-1 remaining tasks $Y\setminus\{y_{c}\}$ is explicitly excluded from the training process.
@@
Within the prior branches, $\varepsilon_{i}$ processes an identical input image, generating a corresponding spatial embedding $e_{i}$ from prior branch and $e_{c}$ from canonical task branch. Subsequently, global average pooling (GAP) is applied to $e_{i}$ to derive a prior tokens $T=\{t_{1},t_{2},...,t_{n-1},t_{c}\}$. These prior tokens represent disparate embeddings, each encapsulating distinct characteristics that diverge from those of the canonical task. The prior token derived from the canonical task branch is denoted as $t_{c}$. To accommodate variations in prior token dimensionality arising from differing encoder capacities, zero-padding is employed to mask any remaining space up to the uniform size. Zero-padding enforce dimensional uniformity by masking vacant spatial space with zeros, thereby obviating superfluous linear transform computations.
3.2 Learning ensemble token
@@
We introduce the ET-Fuser architecture to enable the learning of ensemble tokens. ET-Fuser functions to adapt a ensemble token, drawing inspiration from class token within attention mechanisms and object tokens in DETR [2], to a designated canonical task through the utilization of prior tokens. Leveraging a self-attention mechanism, ET-Fuser facilitates the learning of both ensemble token $t_{k}$ in ET-Fuser and canonical prior $t_{c}$ derived from canonical encoder $e_{c}$. By focusing on the canonical task, it effectively reduces redundancy among the remaining prior tokens, which represent task-optimized embeddings from pre-trained tasks, thereby generating plenty representations $\hat{t}_{k}$ and $\hat{t}_{c}$. Distinct from ensemble and ControlNet architectures, the ET-Fuser architecture eliminates the requirement for supplementary priors during inference, thereby enabling computationally efficient inference based solely on the learned ensemble token.
@@
In Fig. 4, we initialize the $t_{k}$, a learnable query feature, and subsequently concatenate it with the $T$. Drawing inspiration from MaskFormer [4] and dropout [36], random subsampling within the feature representation not only mitigates overfitting but also rectifies global representations and potential inconsistencies in the token’s forward pass between training and inference phases, specifically within ET-Fuser. We introduce a stochastic drop mask $m$ denoted as:
@@
$T_{masked}=\{t_{i}\in T\mid m_{i}=1\},$ (2)
which randomly subsamples the $T$ over threshold $\theta$. The $T$ are subjected to drop mask and consequently excluded from further processing. During the construction of the $m_{i}$, the $t_{c}$ is invariably passed through the subsampling, ensuring inference consistency and conferring a higher magnitude relative to other prior tokens, mirroring the treatment of the ensemble token itself.
We employed multi-head self-attention (MHSA) for learning ensemble tokens. The subsampled token set $T_{{masked}}$ in Eq. 2 is presented as input to the conventional attention mechanism:
@@
Figure 4: Detailed structure of ET-Fuser. The ET-Fuser is trained on a restricted information $T_{\text{mask}}$ to enrich the $e_{c}$ in Fig. 3. During the training phase, the MHSA layer processes $T_{\text{mask}}$, which has been subjected to subsampling via the drop mask. During inference, the MHSA layer only processes $\{t_{k},t_{c}\}$, and the canonical mask is omitted.
In application of self-attention, positional embeddings are applied to $T_{\text{masked}}$ based on the primitive task position $i$. $\hat{T}$ generated by the attention layer undergoes subsampling to ${\hat{t}_{k},\hat{t}_{c}}$ via a canonical mask. Significantly analogous to the stochastic drop mask, the canonical mask serves to mitigate potential inconsistencies between the training and inference phases. In subsequent, a feed-forward network and normalization layer facilitate a seamless transition of the plenty representations ${\hat{t}_{k},\hat{t}_{c}}$ to the decoder pathway.

### v2

@@ Abstract
Facial analysis exhibits task-specific feature variations. While Convolutional Neural Networks (CNNs) have enabled the fine-grained representation of spatial information, Vision Transformers (ViTs) have facilitated the representation of semantic information at the patch level. While advances in backbone architectures have improved over the past decade, combining high-fidelity models often incurs computational costs on feature representation perspective. In this work, we introduce KT-Adapter, a novel methodology for learning knowledge token which enables the integration of high-fidelity feature representation in computationally efficient manner. Specifically, we propose a robust prior unification learning method that generates a knowledge token within a self-attention mechanism, sharing the mutual information across the pre-trained encoders. This knowledge token approach offers high efficiency with negligible computational cost. Our results show improved performance across facial analysis, with statistically significant enhancements observed in the feature representations.
@@ 1 Introduction
Facial analysis exhibits task-specific feature variations; conversely, a unified feature representation derived from diverse features yields robust performance across facial analysis tasks. Deep learning neural networks represent spatial and semantic features from a single facial image. The Convolutional Neural Networks (CNNs) enable spatial information representation, and more recently, the Vision Transformers (ViTs) facilitate the semantic information representation at the patch level. Notably, the transformer’s attention mechanism enables robust representation even when inputs are scattered across disparate spaces or modalities [13, 40].
@@
Figure 1: The key concept of our proposed approach. Our method leverages pre-trained priors from a frozen encoder and unifies them into a learnable token, called knowledge token.
@@
The emergence of general models such as DINOv2 [25] and SAM [14] capable of handling tasks has significantly advanced interpretability within the computer vision. These general models have demonstrated significant achievements in image classification, segmentation, and detection, and as well as handling multi-tasks within a single encoder. Furthermore, linear probing of pre-trained models with unified feature representation has shown promising results [10, 16, 27, 12]. In the areas of multi-modality, CLIP [29, 32] and attention mechanisms [38, 5] have demonstrated promising results, confirming that unified feature representation significantly enhances model’s interpretability.
@@
As noted, the capacity of single and multi-task learning in conventional methodologies has stepped up the visual interpretability. However, there remains a paucity of research that preserves the unified feature representation or encoder’s priors on single task learning during the training phase. Although task-oriented fine-tuning can be achieved through ensemble learning [8, 21, 22], there is still a challenge in that the computational cost escalates proportionally to both the number of tasks and the encoder model’s capacity. In this research, we introduce a methodology for learning knowledge tokens by leveraging the attention mechanisms based on the task priors of pre-trained models in facial analysis tasks. The principal contributions of this study are shown in  1 and as follows:
  • We propose KT-Adapter, a module for adapting and learning query priors based on self-attention mechanism, sharing the mutual information across the prior tasks.
@@
  • The knowledge token within the KT-Adapter denotes a highly efficient method and requires negligible computational cost. Furthermore, the KT-Adapter’s flexible architecture allows for straightforward adaptation and integration across domains and tasks
@@
  • We present an efficient task-specific adaptation without training from scratch, unaffected by data limitations and hardware cost.
@@ 2 Related works
Figure 2: Knowledge learning typically train (red) pre-trained (blue) priors within a fusion module, while ControlNet incorporates an auxiliary lightweight (dashed-line) encoder. In contrast, KT-Adapter implicitly (grey) learns encoder priors within its fusion module, thereby reducing computational cost dramatically.
@@ 2.1 Leveraging Pre-trained Priors
Ensemble learning. The capacity to train specialized models for individual tasks has enabled ensemble methods to achieve robust performance through the exploitation of task-specific priors. These methods range from classical techniques like bagging, boosting, and stacking, to more sophisticated fusion approaches such as majority voting and deep learning-based methodologies [28, 2, 17, 20]. Nevertheless, a fundamental limitation of ensemble methodologies resides in their computational cost, which is directly proportional to both model capacity and ensemble cardinality. This computational cost can preclude their applicability in real-time facial analysis scenarios as shown in  2. Although recent studies [42, 39] have explored strategies to mitigate the computational cost of model cardinality, the challenge of developing training frameworks that are simultaneously efficient, learnable, and adaptable across high feature spaces remains unresolved.
Low rank adaptation (LoRA). The LoRA [11] was introduced to enable parameter-efficient fine-tuning (PEFT), primarily to reduce the computational cost and hardware demands of training parameter-heavy models. Within the computer vision, LoRA has also proven valuable as a PEFT strategy, especially for diffusion models [9, 7]. Following LoRA, ControlNet-based [44] adaptation of foundation models has demonstrated significant efficacy in domain and task adaptation even on multi-modality face analysis tasks [15, 36]. While ControlNet has established its capacity to preserve pre-trained priors within task-specific models, the potential for ensembling encoders to achieve controllable training warrants further investigation.
@@ 2.2 Feature Tokenization on Facial Analysis
The attention mechanism [5] and CLIP [26] have become a cornerstones within the computer vision, effectively capturing both the semantic tokens and the spatial hierarchies within single and multi-modal data. In facial analysis, the HOTformer [34] introduced a novel approach employing lightweight CNNs guided by holistic and atomic tokens. This hybrid token of HOTformer effectively integrates facial semantics and outperforms the accuracy on face recognition. The preliminary work [45] proposed TokenFace, a transformer-based 3D face reconstruction model with a learnable facial component tokens encompassing shape, expressions, jaw pose, camera pose, texture, and lighting. This multi-modality approach effectively integrates the interdependencies between facial characteristics and image token embeddings derived from ViT within the single task. The POSTER++ [19] advanced the state-of-the-art in facial emotion recognition by replacing traditional pyramid architectures with a more efficient design, achieving performance with minimal computational cost. This method leverages cross-attention to fuse landmark priors and image features. The FaceXFormer [24] fuses the hierarchical representation extracted from CNN blocks. Its task-specific tokenization scheme facilitates multi-task learning within a single encoder architecture, effectively utilizing both self-attention and cross-attention mechanisms.
@@
Figure 3: Training phase of proposed method. Within the designated blue block, all encoders $E$ are maintained as frozen layers, indicating that they are not subject to parameter updates during training. Both the KT-Adapter and the canonical branch decoder $d_{c}$ represent the trainable layer within the architecture. In the green block, decoders belonging to the set $D\setminus\{d_{c}\}$, which correspond to other tasks distinct from the canonical task, are excluded from the forward pass graph. Furthermore, The existence of the ground truth set $Y\setminus\{y_{c}\}$, denoted as dashed line, is contingent upon the dataset, exhibiting potential variability in its presence; exceptionally, the $y_{c}$ associated with the canonical branch is an indispensable element.
While attention mechanisms have demonstrated efficacy in capturing mutual information at both coarse and fine-grained levels, and while facial analysis benefits from the flexibility of determining between CNN- and ViT-based architecture contingent on task-specific requirements, current methodologies employ a single learnable encoder for mutual information learning. This architectural characteristic can induce performance bottlenecks, particularly when constrained by encoder’s capacity. So as in leveraging the pre-trained priors, ensemble learning often achieve state-of-the-art results, but their computational expense can preclude their practical application on facial analysis. Although LoRA-based architectures transform embedding representations via auxiliary layers, the use of a single, lightweight encoder prove insufficient to fully capture the overfull information encoded within each pre-trained priors. Our KT-Adapter addresses both the computational cost and the issue of prior’s redundancies through the knowledge token.
@@ 3 Methodology
Inspired by the object token in DETR [1] and mask token in MaskFormer [3, 4], we propose learnable knowledge tokens corresponding to learnable query features in preliminaries. DETR introduced an anchor-free semantic representation leveraging object tokens. Subsequently, Mask2Former utilized masked attention to incorporate both semantic and spatial information into learnable query features. Motivated by the concept of learnable query embeddings, we introduce a knowledge token designed to capture diverse embeddings derived from the priors of a pre-trained model. While conceptually similar to a module that fuses information acquired from an ensemble, this knowledge token is specifically tailored to the canonical task, selectively attending to and learning pertinent semantic information.
3.1 Acquiring Prior Tokens
@@
In the preliminary stage, a set of N pre-trained encoders $E=\{\varepsilon_{1},\varepsilon_{2},...,\varepsilon_{n-1},\varepsilon_{c}\}$ corresponding to the expertise of each task is prepared. These encoders could be employed either CNN or ViT architectures, or even a hybrid thereof dictated by task-specific priors. Subsequent to the preparation of $E$, a canonical task branch is selected for task-specific training. As illustrated in  3, face landmark detection serves as the canonical task in the presented example. During training stage, supervision learning is exclusively derived from the ground truth $Y=\{y_{1},y_{2},...,y_{n-1},y_{c}\}$ corresponding to the designated canonical task; ground truth of the N-1 remaining tasks $Y\setminus\{y_{c}\}$ is explicitly excluded from the training process.
@@
Within the prior branches, $\varepsilon_{i}$ processes an identical input image, generating a corresponding spatial embedding $e_{i}$ from prior branch and $e_{c}$ from canonical task branch. Subsequently, global average pooling (GAP) is applied to $e_{i}$ to derive a prior tokens $T=\{t_{1},t_{2},...,t_{n-1},t_{c}\}$. These prior tokens represent disparate embeddings, each encapsulating distinct characteristics that diverge from those of the canonical task. The prior token derived from the canonical task branch is denoted as $t_{c}$. To accommodate variations in prior token dimensionality arising from differing encoder capacities, zero-padding is employed to mask any remaining space up to the uniform size. Zero-padding enforces dimensional uniformity by masking unused spatial positions with zeros, thereby obviating superfluous linear transform computations.
Figure 4: Detailed structure of KT-Adapter. The KT-Adapter is trained on a restricted information $T_{\text{mask}}$ to enrich the $e_{c}$ in  3. During the training phase, the MHSA layer processes $T_{\text{mask}}$, which has been subjected to subsampling via the drop mask. During inference, the MHSA layer only processes $\{t_{k},t_{c}\}$, and the canonical mask is omitted.
3.2 KT-Adapter
@@
We introduce the KT-Adapter architecture to enable the learning of knowledge tokens. KT-Adapter functions to adapt a knowledge token, drawing inspiration from class token within attention mechanisms and object tokens in DETR [1], to a designated canonical task through the utilization of prior tokens. Leveraging a self-attention mechanism, KT-Adapter facilitates the learning of both knowledge token $t_{k}$ in KT-Adapter and canonical prior $t_{c}$ derived from canonical encoder $e_{c}$. By focusing on the canonical task, it effectively reduces redundancy among the remaining prior tokens, which represent task-optimized embeddings from pre-trained tasks, thereby generating representations $\hat{t}_{k}$ and $\hat{t}_{c}$. Distinct from ensemble and ControlNet architectures, the KT-Adapter architecture eliminates the requirement for supplementary priors during inference, thereby enabling computationally efficient inference based solely on the learned knowledge token.
3.3 Drop mask and Canonical mask
@@
In  4, we initialize the $t_{k}$, a learnable query feature, and subsequently concatenate it with the $T$. Drawing inspiration from MaskFormer [3] and dropout [33], random subsampling within the feature representation not only mitigates overfitting but also rectifies global representations and potential misalignment in the token’s forward pass between training and inference phases, specifically within KT-Adapter. We introduce a stochastic drop mask $m$ denoted as:
@@
$T_{mask}=\{t_{i}\in T\mid m_{i}=1\},$ (2)
which randomly subsamples the $T$ over threshold $\theta$. The $T$ are subjected to drop mask and consequently excluded from further processing. During the construction of the $m_{i}$, the $t_{c}$ is invariably passed through the subsampling, ensuring inference consistency and conferring a higher magnitude relative to other prior tokens, mirroring the treatment of the knowledge token itself.
We employed multi-head self-attention (MHSA) for learning knowledge tokens. The subsampled token set $T_{{masked}}$ in  2 is presented as input to the conventional attention mechanism:
@@
In application of self-attention, positional embeddings are applied to $T_{\text{masked}}$ based on the primitive task position $i$. The $\hat{T}$ generated by the attention layer undergoes subsampling to $\hat{t}_{k},\hat{t}_{c}$} via a canonical mask. Significantly analogous to the stochastic drop mask, the canonical mask serves to mitigate potential inconsistencies between the training and inference phases. In subsequent, a feed-forward network and normalization layer facilitate a seamless transition of the representations $\hat{t}_{k},\hat{t}_{c}$} to the decoder pathway.